OpenStudy (anonymous):

how would you use the the following matrix l x y z l l u v w l = 4 l 1 2 3 l what would the determinant of the bottom matrix l x y z l l -4 -8 -12 l l u v w l

OpenStudy (anonymous):

Whenever we swap two rows of a matrix, the sign of the determinant changes: If \[\det \left[\begin{matrix}a & b \\ c & d\end{matrix}\right]= e\] Then \[\det \left[\begin{matrix}c & d \\ a & b\end{matrix}\right]= -e\] Also, when we do a scalar multiplication of a row, whatever the value of that scalar multiplications is also done on the determinant. \[\det \left[\begin{matrix}a & b \\ c & d\end{matrix}\right]= e\] Then \[\det \left[\begin{matrix}a & b \\ 2c & 2d\end{matrix}\right]= 2e\]

OpenStudy (anonymous):

i multiplied the it just by -4 didnt know the sign changes! Thanks alot!